A splash is heard after 2.05 secs after a stone was dropped into a well of depth 19.6 m. The velocity of sound is??

Grade:11

1 Answers

Vikas TU
7 years ago
Actual  time to raach the bottom=>
19.6 = 0 + 0.5gt^2
t = root(3.92) = > 1.9798 seconds.
Delta t = > 2.05 – 1.9798 = > 0.07010 seconds.
v = 19.6/0.07010 = > 279.6 m/s.